14.34 g of an oxide of lead was heated in a stream of dry hydrogen, 12.42 g of lead was obtained. Relative atomic masses of oxygen and lead are 16 and 207 respectively.
1. How many grams of oxygen were presthe oxide?
2. Calculate the number of moles of oxygen atoms present.
1
Expert's answer
2017-02-01T13:01:12-0500
PbO + H2 = Pb + H2O
We use the law of equivalence
m(O) / Мeq (О) = m(Pb) / Мeq (Pb)
In these compounds with one mole of lead atoms connected one mole of oxygen atoms

m(O) = m(Pb) x Мeq (О) / Мeq (Pb)
m(O) = 16 g/mol x 12.42 g / 207 g/mol = 0.96 g
Determine the quantity of the substance in oxygen 0.96 g
n(O) = m (O) / M (O) = 0.96 / 16 = 0.06 mol